find wave velocity when in shm when y=0.25cos (125t- 0.25x)
1
Expert's answer
2020-06-17T09:10:07-0400

We are assuming all the quantities are in S.I.

Wave equation is

"y=0.25\cos(125t-0.25x)=0.25\sin(0.25x-125t+\pi)"

Now,

"k=0.25\implies \lambda=\frac{2\pi}{0.25}=8\pi"

And,

"\frac{2\pi v_w}{\lambda}=125\implies v_w=125\cdot4=500m/s"